The sum of 2 numbers is 40, if 2 is added to the larger the result is equal to twice the smaller number. What are the two numbers?.
Answer by rwm(914)

a+b=40
a+2=2b
a=2b-2
2b-2+b=40
3b=42
b=14
a=26